### v1

# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

Abstract
This paper introduces a framework that integrates image compression and super-resolution using a unified representation called Generalizable Factor Fields. Traditional approaches to image compression often lead to significant information loss, particularly at low bitrates, which blocks image usability for both human viewing and computer vision applications. Super-resolution techniques strive to reconstruct high-quality images from these degraded compressions but face challenges when starting with heavily compressed sources. Our approach leverages a coefficient-basis structure within the Factor Fields, optimizing the joint task of compressing multiple images and enhancing their resolution in a single cohesive process. By encoding and transmitting images using shared basis functions, our method significantly reduces information loss and boosts efficiency in handling large image datasets. We demonstrate superior performance in both super-resolution and compression tasks compared to existing methods, offering improved image reconstruction quality and more efficient low-bitrate operation.
1 Introduction
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches flickr2k, SRCNNlaid the foundation for SISR, which was later enhanced with GAN-based methods srgan for improved perceptual realism. Follow-up Transformer-based networks IPT address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements swin2sr, DAT, RGT, which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards VVC-21.2, jpeg, jpeg2000 lay the groundwork. The emerging learned image compression models imc_cnn1, ballé2018variational, guo-hua2023evc, imc_lic_tcm, minnen2018joint, cheng2020learned, compression algorithms of which mostly follow the pixel-space transform coding  imc_cnn1, Goyal2001paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling tensorf, mueller2022instantngp, eg3d, kplanes, cao2023hexplane, factor-field, gao2023strivec, we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
Finally, we propose merging the bases by introducing an additional compression branch, consolidating multiple bases into one alongside multi-image transmission. This approach leverages the mutual information across multiple images, reducing the need for redundant transmissions and refining the basis structure. The main contributions of this paper are summarized as follows:
  • We propose Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
  • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
  • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
2 Related Works
Super-Resolution (SR).
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies SRCNN, flickr2k, zhou2020cross, shufflemixer, VDSR were initially introduced with modeling techniques such as residual learning ledig2017photo, liu2020residual, tong2017image, zhang2018residual, RCAN, drct, EDSR , or recursive learning RGT, tai2017recursive. Besides, subsequent research also sheds light on GAN-base methods srgan, ledig2017photo, esrgan, real_esrgan to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks IPT, EDT. Afterward, SwinIR swinir is proposed to combine spatial locality and non-local information by Swin Transformer liu2021swin with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework swin2sr, zhu2023attention, ATD to reach better image quality as well as solve information bottleneck drct. Hybrid approaches CRAFT (CRAFT) merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT (DAT) and HAT (chen2023hat) integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT (RGT) introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
Image Compression (IC).
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG jpeg and JPEG-2000 jpeg2000. Early CNN-based approaches imc_cnn1, imc_cnn2, imc_cnn3 have been surpassed by transformer-based models imc_trans1, imc_trans2, imc_trans3, duan2023cross-component, which leverage spatio-channel attention for better performance. GAN-based methods imc_gan1, imc_gan2, imc_gan3, imc_gan4 have further contributed to real-time adaptive compression. Recently, ELIC imc_elic introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM imc_lic_tcm integrated transformers and CNNs to capture both global and local image features. The eContextformer koyuncu2024efficient introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer li2024groupedmixer proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion song2024wavelet-compression introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach. We propose Factorized Fields, a unified framework that models visual and structural features, offering a more comprehensive solution to enhance performance in both tasks.
3 Methods
In this section, we first briefly introduce the background of the factor fields factor-field, their properties, and also the preliminary of Learned Image Compression in Sec. 3.1. Inspired by factor fields, we next explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in Sec. 3.2. We then describe how to adapt the formulation to Super-Resolution in Sec. 3.3. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in Sec. 3.4.
Figure 1: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly.
3.1 Preliminaries
Factor Fields.
The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently tensorf, gao2023strivec, mueller2022instantngp, kplanes, cao2023hexplane, eg3d, while factor-field first proposed a unified representation called factor fields. For a 1D signal $s(x)$, the factor fields formulation is:
$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (1)
where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation. The coefficient and basis are queried by sampling with the coordinates $x$. This generalizes to Q-dimensional signals:
$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (2)
where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
Learned Image Compression.
Following (channel-autoregressive, ballé2018variational), a learned image compression model with a channel-wise entropy model can be formulated as:
$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (4)
$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z)$
$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},\ 0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}}))$
$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}).$
The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is quantized to $\hat{z}$, which is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow imc_lic_tcm, modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
Figure 2: The overall pipeline of image super-resolution with our Factorized Fields. Given a low-resolution image $I_{LR}$, we first extract coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficient and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
3.2 Formulation of Factorized Fields
As discussed in Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information fuoli2021fourier, waveletsr to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields kplanes, factor-field, tensorf, mueller2022instantngp, cao2023hexplane, eg3d, we primarily built our Factorized Fields framework on factor fields factor-field from Eq. 3:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
Note that $\hat{I}$ denotes the approximated images, and $x\in R^{2}$ are the pixel coordinates.
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated. Finally, as in factor-field, of all the tested coordinate transformation $\gamma$ in Eq. 3 sawtooth $\gamma(x)=x\bmod{k},k\in\mathbb{R}$ performs mostly the best in image regression tasks. We can easily observe that such transformation implicitly captures patch-like frequency information as shown in Fig. 1a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
However, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R.$ (7)
Here, inspired by the Fourier Series, with a scalar $\alpha$ which can be viewed as increasing to sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
Figure 3: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec.3.3 as information-recovery prior, detailed in Sec. 3.4.1. (c) Expanding on (b), Sec. 3.4.2 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.3 Super-Resolution with Factorized Fields
We represent a super-resolved image using our Factorized Fields, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
where $c^{\text{LR}}_{i}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(x)$ and $b^{\text{LR}}_{i}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(\gamma_{i}(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in Fig. 2, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using Eq. 8, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods chen2023activating, DAT, ATD as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks liu2022swinv2 with a series of downsampling operations. We use a dilation-like downsampling technique (Fig. 1(c)) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.4 Image Compression with Factorized Fields and Super-Resolution
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in Sec. 1,  2, and 3.2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models koyuncu2024efficient, li2024groupedmixer, or decompression modules song2024wavelet-compression, duan2023cross-component, while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in Eq. 7. In addition, with our trained SR model described in Sec. 3.3, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
3.4.1 Single Image Compression
The overall pipeline is shown in Fig. 3(b). To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow imc_lic_tcm, with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined in Eq. 5.
3.4.2 Multi-Image Compression
For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure. These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise. Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),\ 0\leq h<H_{b_{i}},\ 0\leq w<W_{b_{i}}.$ (9)
Here, $F_{\text{merge}}$ is a standard transformer, similar in architecture to that described in (dinov2). We treat the bases as tokens and prepend the sequence with a $CLS$ token, concluding the sequence to form the final merged basis.
Since compression induces error, according to Fig. 3b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error. To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients. as illustrated in Fig. 3c. This way, we can enjoy less information loss with basis while maintaining low bit rates.
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining chen2023activating. Please refer to quantitative results in Sec. 4.1 for details.
Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
[…]
SwinIR (swinir)(ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD (zhang2024transcending)(CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT (DAT))(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT (RGT)(ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT (chen2023hat)(CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L (chen2023hat)(CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
ATD-F (Ours) 45.46 149.87 8,674 33.29 0.9082 29.48 0.8017 28.03 0.7539 28.53 0.8487 33.11 0.9335
DAT-F (Ours) 40.00 134.42 6,206 33.45 0.9094 29.60 0.8039 28.13 0.7560 28.75 0.8520 33.23 0.9339
HAT-F (Ours) 45.97 158.79 5,750 33.53 0.9100 29.65 0.8050 28.18 0.7569 28.79 0.8527 33.33 0.9342
HAT-F-ImageNet (Ours) 45.97 158.79 5,750 33.55 0.9102 29.63 0.8049 28.18 0.7569 28.80 0.8529 33.33 0.9342
HAT-L-F (Ours) 66.04 240.03 8,888 33.75 0.9116 29.87 0.8091 28.31 0.7597 29.51 0.8637 33.36 0.9343
HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
[…]
4 Experiments
4.1 Image Super-Resolution
Experimental Setup.
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in (chen2023hat) and (drct), we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from (sam), which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet (imagenet), which is used for chen2023hat and drct pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR (swinir), HAT (chen2023activating), DAT (DAT), and ATD (zhang2024transcending)) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K (div2k) + Flickr2K (flickr2k)) as the finetuning dataset following (swin2sr, chen2023hat) for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
Quantitative Results.
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 (set5), Set14 (set14), BSD100 (bsd100), Urban100 (urban100), and Manga109 (manga109). For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD ATD, DAT DAT, and HAT chen2023activating—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only chen2023activating provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in IPT, chen2023activating. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods jpeg, the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in Fig. 2. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison.
We provide the visual comparison in Fig. 4. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
Tot Enc $\downarrow$ Tot Dec $\downarrow$
VTM 0.00 129.21 0.14 -
Xie (MM 21’) -0.78 2.93 6.00 50.0
Cheng (CVPR 22’) 5.44 1.98 4.69 29.6
STF (CVPR 22’) -4.31 0.14 0.13 99.9
ELIC (CVPR 22’) -7.24 0.07 0.09 36.9
TCM (CVPR 23’) -11.74 0.16 0.15 76.7
TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
{subfigure}
[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption
Figure 5: Kodak
Figure 6: CLIC
Figure 7: Tecnick
Figure 8: Performance (RD-Curve) evaluation on image compression using different datasets.
4.2 Single- and Multi-Image Compressions
Experimental Setup.
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig.3 from pre-trained imc_lic_tcm and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer (loshchilov2017decoupled) with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 3(c), we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
Rate-Distortion Performance Comparison.
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including (liu2023learned), (Chen2022), zou2022the, xie2021enhanced, cheng2020learned, ballé2018variational, li2024groupedmixer, jiang2023mlic, channel-autoregressive, bellard_bpg, qi2023hyperthumbnail, and imc_elic. The classical image compression codec, VVC (vtm), is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
In Tab. 2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods. The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$. The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
Our analysis reveals several significant advantages of the FIPER framework in image compression tasks. The approach demonstrates substantial improvements in compression efficiency across various benchmark datasets, including Kodak, CLIC, and Tecnick, indicating its broad applicability. Notably, the multi-image compression strategy shows particularly promising results for larger image sets, suggesting scalability benefits. Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
Table 3: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 8
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline (factor-field) 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
4.3 Ablation Studies
Effectiveness of Factorized Fields Design.
We conduct experiments to verify our modification of Factorized Fields, modified from Eq. 3 to Eq. 7. The quantitative performance reported on single-image regression is shown in Tab. 3, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 4, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT chen2023activating in the pipeline illustrated in Fig. 3.b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
Table 4: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
Method Kodak CLIC Tecnick
$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
[…]
TCMimc_lic_tcm 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
5 Conclusion
We proposed Factorized Fields, a representation that models implicit structures and patterns by decomposing images into multi-frequency components. This approach addresses challenges in Super-Resolution and Image Compression by restoring details and preserving visual fidelity. We integrate SR priors with Image Compression for improved information recovery and introduce a basis merging technique for enhanced rendering quality across multiple images. Experiments demonstrate state-of-the-art performance in both SISR and Image Compression benchmarks, addressing limitations of previous methods.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
[…]
    means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
[…]
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
[…]
  4. Justification: Our abstract and introduction accurately reflect the paper’s contributions and scope by highlighting our novel contribution: we employ the factor fields for enhance the performance in super-resolution and multi-image compression.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
[…]
  9. Justification: Our paper explicitly discusses the limitations of our approach, including the long training and inference time.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
[…]
  14. Justification: Our paper primarily focuses on empirical research and does not introduce new theoretical results. As such, there are no theoretical assumptions or proofs discussed in the study.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
[…]
  19. Justification: Our paper thoroughly details the methodologies, parameter settings, and experimental setups, providing all necessary information for replicating the main experimental results. This comprehensive documentation supports the validity of our primary claims and conclusions, ensuring clarity and reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
[…]
  24. Justification: We will make a full open access to both the code and the data used in our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
[…]
  29. Justification: Our paper meticulously details the experimental setup, including data splits for training and testing, hyperparameter configurations, and the types of optimizers utilized. These elements are crucial for a thorough understanding of the experimental procedures and the interpretation of results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
[…]
  34. Justification: Due to the limited scope of our experiments, which encompass only two trials yielding similar results, our paper does not include detailed error bars or extensive statistical analysis. We recognize the potential for variability in these findings and suggest that additional experiments be conducted under varied conditions to validate and expand upon our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
[…]
  39. Justification: We use GPU GTX 4090 as our computational resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
[…]
  44. Justification: We carefully follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
[…]
  49. Justification: In the conclusion of our paper, we discuss the societal impacts directly related to our research on super resolution and image compression.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
[…]
  54. Justification: Our paper does not include the release of pre-trained models, new datasets, or any other resources that typically necessitate specific safeguards to prevent misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
[…]
  59. Justification: We have diligently credited the creators and original owners of all assets used in our research, including code, data, and models, ensuring that each source is accurately cited within our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
[…]
  64. Justification:: Yes, any new assets introduced in the paper will be well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
[…]
  69. Justification: Our research does not involve crowdsourcing experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
[…]
  74. Justification: Our research does not involve human subjects, and therefore, there were no study participants exposed to any risks.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.