### v2

# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

Abstract
In this work, we propose a unified representation for Super-Resolution (SR) and Image Compression, termed Factorized Fields, motivated by the shared principles between these two tasks. Both SISR and Image Compression require recovering and preserving fine image details—whether by enhancing resolution or reconstructing compressed data. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition to explicitly capture multi-scale visual features and structural components in images, addressing the core challenges of both tasks. We first derive our SR model, which includes a Coefficient Backbone and Basis Swin Transformer for generalizable Factorized Fields. Then, to further unify these two tasks, we leverage the strong information-recovery capabilities of the trained SR modules as priors in the compression pipeline, improving both compression efficiency and detail reconstruction. Additionally, we introduce a merged-basis compression branch that consolidates shared structures, further optimizing the compression process. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
1 Introduction
Figure 1: The Proposed Factorized Fields. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, especially useful for both Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=sin$.
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [46, 21]laid the foundation for SISR, which was later enhanced with GAN-based methods [39] for improved perceptual realism. Follow-up Transformer-based networks [12] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [19, 15, 16], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [33, 69, 66] lay the groundwork. The emerging learned image compression models [7, 2, 29, 48, 58, 17], compression algorithms of which mostly follow the pixel-space transform coding  [7, 27]paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [9, 59, 8, 24, 6, 10, 26], we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
Finally, we propose merging the bases by introducing an additional compression branch, consolidating multiple bases into one alongside multi-image transmission. This approach leverages the mutual information across multiple images, reducing the need for redundant transmissions and refining the basis structure. The main contributions of this paper are summarized as follows:
  • We propose Generalized Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
  • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
  • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
2 Related Works
Super-Resolution (SR).
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies [21, 46, 79, 64, 34] were initially introduced with modeling techniques such as residual learning [40, 47, 68, 78, 77, 18, 45] , or recursive learning [16, 65]. Besides, subsequent research also sheds light on GAN-base methods [39, 40, 70, 71] to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks [12, 43]. Afterward, SwinIR [44] is proposed to combine spatial locality and non-local information by Swin Transformer [50] with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework [19, 80, 75] to reach better image quality as well as solve information bottleneck [18]. Hybrid approaches CRAFT [41] merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT [15] and HAT [13] integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT [16] introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
Image Compression (IC).
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG [69] and JPEG-2000 [66]. Early CNN-based approaches [7, 3, 28] have been surpassed by transformer-based models [37, 81, 72, 22], which leverage spatio-channel attention for better performance. GAN-based methods [53, 54, 23, 62] have further contributed to real-time adaptive compression. Recently, ELIC [30] introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM [48] integrated transformers and CNNs to capture both global and local image features. The eContextformer [38] introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer [42] proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion [63] introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach. We propose Factorized Fields, a unified framework that models visual and structural features, offering a more comprehensive solution to enhance performance in both tasks.
3 Methods
In this section, we present the details of our proposed methods. Inspired by factor fields, which can be formulated by
$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
and whose details can be found in supplementary materials , we first explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in  Sec. 3.1. We then describe how to adapt the formulation to Super-Resolution in  Sec. 3.2. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in  Sec. 3.3. Note that the background of learned image compression is described in supplementary materials.
Figure 2: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
Figure 3: The overall pipeline of image super-resolution with our Factorized Fields. Given a low-resolution image $I_{LR}$, we first extract coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficient and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 4, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
3.1 Formulation of Factorized Fields
As discussed in  Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information [25, 36] to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields [24, 10, 9, 59, 6, 8], we primarily built our Factorized Fields framework on factor fields [10] from  Eq. 1:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
Note that $\hat{I}$ denotes the approximated images, and $x\in R^{2}$ are the pixel coordinates.
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated.
A key insight of our work is the use of sawtooth coordinate transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, which performs best among tested transformations in image regression tasks. The sawtooth transformation maps input coordinates to a periodic space, implicitly capturing patch-like frequency information as shown in Fig. 2a. However, traditional downsampling in neural networks fails to preserve this periodic structure. To address this, we introduce sawtooth-aware downsampling as in Fig. 2c with our Basis Swin Transformer. This rearrangement allows the network to learn correlations between the periodic features in Sawtooth sampling.
Moreover, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (3)
where $\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R$. Here, inspired by the Fourier Series, with a scalar $\alpha$ which corresponds to an increase in sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
Refer to captionRefer to caption
Figure 4: Comparison of Different Representation This figure illustrates the performance comparison on a single-image regression task using three different representations: vanilla factor fields in Eq. 1, our Factorized Fields (Eq. 3), and Plain Fields, which use simple channel-wise tensor concatenation instead of coefficient-basis multiplication. The left panel shows the MAE on the magnitude of different frequency components obtained from the Fourier Transform, while the right panel shows PSNR along training. Although the basis-coefficient decomposition method in [10] yields improvements, it still struggles with slow convergence and inconsistencies between low and high-frequency components. In contrast, by introducing explicit formulations for $\alpha$ and $\psi$, our method aligns different frequency components effectively and achieves superior convergence speed. This advantage becomes especially crucial when the task extends to generalizable applications, such as super-resolution or image compression.
Our Factorized Fields enhance the original Factor Fields [10] in three key ways: (1) non-linear transformation through $\psi\in\{\text{sin},\text{cos}\}$ enables better modeling of complex image patterns, (2) multi-frequency modulation using $\alpha_{j}$ allows explicit control over frequency components, and (3) sawtooth-aware downsampling preserves periodic structure during feature processing. As shown in Fig. 4, our refined representation contributes to large margin of improvement in convergence and performance, demonstrating the modification is the key to superior information modeling.
Figure 5: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec. 3.2 as information-recovery prior, detailed in Sec. 3.3.1. (c) Expanding on (b),  Sec. 3.3.2 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.2 Super-Resolution with Factorized Fields
We represent a super-resolved image using our Factorized Fields, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (4)
where $c^{\text{LR}}_{i}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(x)$ and $b^{\text{LR}}_{i}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(\gamma_{i}(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 4, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods [14, 15, 75] as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks [51] with a series of downsampling operations. We use a dilation-like downsampling technique( Fig. 2c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.3 Image Compression with Factorized Fields and Super-Resolution
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1,  Sec. 2, and Sec. 3.1, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [38, 42], or decompression modules [63, 22], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in  Eq. 3. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
3.3.1 Single Image Compression
The overall pipeline is shown in  Fig. 5b. To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [48], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
3.3.2 Multi-Image Compression
For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure. These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise. Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (6)
[…]
Since compression induces error, according to  Fig. 5b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error. To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients. as illustrated in Fig. 5c. This way, we can enjoy less information loss with basis while maintaining low bit rates.
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining [14]. Please refer to quantitative results in  Sec. 4.1 for details.
Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
[…]
SwinIR [44](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD [76](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT [15])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT [16](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT [13](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L [13](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
ATD-F (Ours) 45.46 149.87 8,674 33.29 0.9082 29.48 0.8017 28.03 0.7539 28.53 0.8487 33.11 0.9335
DAT-F (Ours) 40.00 134.42 6,206 33.45 0.9094 29.60 0.8039 28.13 0.7560 28.75 0.8520 33.23 0.9339
HAT-F (Ours) 45.97 158.79 5,750 33.53 0.9100 29.65 0.8050 28.18 0.7569 28.79 0.8527 33.33 0.9342
HAT-F-ImageNet (Ours) 45.97 158.79 5,750 33.55 0.9102 29.63 0.8049 28.18 0.7569 28.80 0.8529 33.33 0.9342
HAT-L-F (Ours) 66.04 240.03 8,888 33.75 0.9116 29.87 0.8091 28.31 0.7597 29.51 0.8637 33.36 0.9343
HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
[…]
4 Experiments
4.1 Image Super-Resolution
Experimental Setup.
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in [13] and [18], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [35], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [20], which is used for [13] and [18] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [44], HAT [14], DAT [15], and ATD [76]) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K [1] + Flickr2K [46]) as the finetuning dataset following [19, 13] for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
Quantitative Results.
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [5], Set14 [74], BSD100 [55], Urban100 [31], and Manga109 [56]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [75], DAT [15], and HAT [14]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [14] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [12, 14]. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods [69], the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in  Fig. 3. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 3. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison.
We provide the visual comparison in Fig. 6. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Multi-frame SR.
Due to the space limit, we provide additional comparisons on multi-frame SR in the supplementary materials.
(a) Kodak
(b) CLIC
(c) Tecnick
Figure 7: Performance (RD-Curve) evaluation on image compression using different datasets.
Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
Tot Enc $\downarrow$ Tot Dec $\downarrow$
VTM 0.00 129.21 0.14 -
Xie (MM 21’) -0.78 2.93 6.00 50.0
Cheng (CVPR 22’) 5.44 1.98 4.69 29.6
STF (CVPR 22’) -4.31 0.14 0.13 99.9
ELIC (CVPR 22’) -7.24 0.07 0.09 36.9
TCM (CVPR 23’) -11.74 0.16 0.15 76.7
TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
Method Kodak CLIC Tecnick
$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
[…]
TCM[48] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
4.2 Single- and Multi-Image Compressions
Experimental Setup.
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from  Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig. 5 from pre-trained [48] and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer [52] with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 5.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
Rate-Distortion Performance Comparison.
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [49], [11], [82], [73], [17], [2], [42], [32], [57], [4], [61], and [30]. The classical image compression codec, VVC [67], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 7.
In  Tab. 2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods. The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$. The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
Our analysis reveals several significant advantages of the FIPER framework in image compression tasks. The approach demonstrates substantial improvements in compression efficiency across various benchmark datasets, including Kodak, CLIC, and Tecnick, indicating its broad applicability. Notably, the multi-image compression strategy shows particularly promising results for larger image sets, suggesting scalability benefits. Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline [10] 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
Table 4: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 4.
4.3 Ablation Studies
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 3, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT [14] in the pipeline illustrated in Fig. 5b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
Effectiveness of Factorized Fields Design.
We conduct experiments to verify our modification of Factorized Fields, modified from Eq. 1 to Eq. 3. The quantitative performance reported on single-image regression is shown in Tab. 4, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
5 Conclusion
We proposed Factorized Fields, a representation that models implicit structures and patterns by decomposing images into multi-frequency components. This approach addresses challenges in Super-Resolution and Image Compression by restoring details and preserving visual fidelity. We integrate SR priors with Image Compression for improved information recovery and introduce a basis merging technique for enhanced rendering quality across multiple images. Experiments demonstrate state-of-the-art performance in both SISR and Image Compression benchmarks, addressing limitations of previous methods.
Limitations.
While the FIPER framework offers advances in image compression and super-resolution, it requires future work to accommodate computation-limited scenarios such as real-time decoding. Also, Factorized Fields don’t consider semantic information, and we will leave this to future research for more possibilities.