### v3

# FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models

Abstract
In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression. Motivated by the shared principles between these tasks, they require recovering and preserving fine image details—whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks. We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability. In addition, we further optimize the pipelines by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression and super-resolution. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA. Please see our project page at jayisaking.github.io/FIPER.
1 Introduction
Figure 1: The Proposed Factorized Features. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=\text{sin}(\cdot)$.
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [52, 23] laid the foundation for SISR, which was later enhanced with GAN-based methods [40] for improved perceptual realism. Follow-up Transformer-based networks [14] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [21, 17, 18], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR? On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [35, 82, 79] lay the groundwork. The emerging learned image compression models [8, 3, 30, 55, 67, 19], compression algorithms of which mostly follow the pixel-space transform coding  [8, 28] paradigm, then introduce a variety of networks  [83, 55] or difference losses [31] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) They aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 68, 9, 26, 7, 12, 27], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively. This approach explicitly captures multi-scale visual features and repetitive structral components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
Finally, we propose merging the bases for multi-image processing by consolidating multiple bases into one through a network, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
The main contributions of this paper are summarized as follows:
  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in Sec. 3.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
  • We generalize such image descriptor to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
  • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
2 Related Works
Super-Resolution (SR). Image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, which is crucial for computer vision. Early CNN-based methods [23, 52, 95, 76, 36] introduced residual [41, 54, 81, 94, 93, 20, 51] and recursive learning [18, 77]. GAN-based approaches [40, 41, 84, 85] improved perceptual quality but faced spatial locality constraints, leading to Transformer-based SISR models [14, 46]. SwinIR [48] integrated window attention with Swin Transformer [58], inspiring further advancements [21, 96, 91, 20]. Hybrid models like CRAFT [43], DAT [17], and HAT [15] optimize feature aggregation, while RGT [18] enhances spatial details efficiently.
Image Compression (IC). Deep learning surpasses traditional compression methods like JPEG [82] and JPEG-2000 [79], with CNN-based [8, 4, 29], Transformer-based [38, 97, 86, 24], and GAN-based [61, 62, 25, 73] approaches improving performance. ELIC [32] introduced adaptive coding, LIC TCM [55] combined CNNs and Transformers, and eContextformer [39] leveraged spatial-channel attention. GroupedMixer [44] introduced token mixers, while Wavelet Conditional Diffusion [75] balanced perceptual quality and distortion.
Our work diverges from these trends by introducing Factorized Features, a unified framework modeling both visual and structural features to enhance performance across super-resolution and image compression.
3 Methods
In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [57], and Factor Fields [12], and delve into the motivation and how we resolve the weaknesses of these formulations by our Factorized Features in Sec. 3.1. Next, we discuss how to integrate such strategy into learned settings of Super-Resolution in Sec. 3.2 and Image Compression Sec. 3.3. Note that the background of learned image compression is provided in supplementary materials.
Figure 2: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
Figure 3: The overall pipeline of image super-resolution with our Factorized Features. Given a low-resolution image $I_{LR}$, we first extract the coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficients and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, and $W$ denote the scale factor, height, and width, respectively.
3.1 Formulation of Factorized Features
Fourier Transform has been long used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions. In practice, it often takes the form of Discrete Cosine Transform for real-valued 2D discrete signal $f$ with height $H$ and width $W$:
$f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right]$ (1)
where $S_{u,v}$ is the amplitude.
Without the loss of generality, we consider square images in this work, with edge length (and also the period) denoted as $T$:
$f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right]$ (2)
where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift. In image processing tasks, the goal is to learn $S$ and $\xi$, then use the formula above to reconstruct high-quality images from degraded inputs. To achieve this, [57] treats $S$ and $\xi$ as learnable parameters, allowing the model to optimize them for better reconstruction. Furthermore, the frequencies $u$ and $v$ are also designed to be promptable by the model, enabling dynamic adaptation to spatial-varying high- and low-frequency components.
More generally, we can formulate it as the multiplication of coefficient $c$ and basis $b$ with simplified notation, and in this work, we use $\mathbf{x}$ to represent the pixel coordinates $(x,y)$:
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),$ (3)
where $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ and $N$ denotes the number of frequency components.
Spatially Variant Coefficient. Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components. Therefore, in this work, we argue that coefficient $c_{i}$ should be spatial-varying, which is different from previous works, i.e.,
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
Learned Non-uniform Basis. Previous works learn basis through frequencies ${u,v}$ and phase shift $\xi$[57, 47, 42]; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details; in other words, since low-frequency components typically require more energy [45], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details, and the predefined nature of these sinusoidal functions further restricts local flexibility, making fine-grained adjustments more challenging. As shown in Fig. 5, higher frequency components converge slower than their counterpart. Hence, in this work, we generate the entire basis map $b_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map and combine them together. In this work, the composition process happens in the end with a projection function $\mathcal{P}$ to form RGB images $\hat{I}$:
$\hat{I}(\mathbf{x})=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl{\}}\Bigl{)}.$ (5)
In contrast to previous works which employ fourier bottlenecks[57], or frequency decomposition blocks[47, 42, 45] among model blocks, here we seek to address a fundamental question: Can a well-learned representation effectively replace the simple feature map output ? That is, our solution is meant to be applied directly on networks to represent any images.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Eq. 4 (b) Eq. 6 (c) Eq. 7 with only $\psi$ (d) Eq. 7 with both $\psi$ and $\alpha$
Figure 4: Visualization of different factorization methods.
Coordinate Transformation. The success of Fourier-based representation originates from its sinusoidal decomposition. By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components. However, the formulation in Eq. 5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in Fig. 4 we witness performance loss due to lack of such guidance. Inspired by position encoding [65, 11, 68, 7, 9] and Fourier feature mapping [78], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, visualized in Fig. 2.a. We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
Multi-frequency Modulation. [72] has shown that neural networks are biased towards learning low-frequency content. Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in Fig. 4:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (7)
where $\psi\in{\sin,\cos}$ and $\alpha\in R$. In this way, each basis $b_{i}$ contributes to both high- and low-frequency components by different values of $\alpha$. Consequently, if a basis is only accurate in the low-frequency domain, the output will contain undesired high-frequency noise.
Alternatively, Eq. 7 can be interpreted as making the spatial coordinates $x,y$ in Eq. 1 learnable instead of the frequency components $u,v$. To validate this, we conduct experiments (Fig. 5) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
To this end, our full Factorized Features representation is formulated in Eq. 7.
Figure 5: In this figure, we validate the claim that fixed sinusoidal functions attend more to low-frequency parts. Learnable basis denotes the Learned Non-uniform basis in Sec. 3.1. And we can see that Learnable Coordinate $x$ has the same effect, verifying our conclusion.
Figure 6: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec. 3.2 as information-recovery prior, detailed in Sec. 3.3 (c) Expanding on (b),  Sec. 3.3.1 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.2 Super-Resolution with Factorized Features
We represent a super-resolved image using our Factorized Features, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
where $c^{\text{LR}}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations. Then, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 8, where $s$ is the scale factor.
We optimize model parameters using $L_{1}$ loss.
To demonstrate the effectiveness of our method, we use existing SR methods [16, 17, 91] as the Coefficient Backbone. For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [59] with a series of special downsampling operations. Specifically, we use a dilation-like downsampling technique( Fig. 2.c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.3 Image Compression with Factorized Features
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1 and  Sec. 2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [39, 44], or decompression modules [75, 24], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through Factorized Features formulation as in  Eq. 7
Priors from SR. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
The overall pipeline is shown in  Fig. 6b. To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [55], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (9)
3.3.1 Multi-Image Compression
For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure. These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise. Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (10)
[…]
Since compression induces error, according to  Fig. 6b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error. To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients. as illustrated in Fig. 6c. This way, we can enjoy less information loss with basis while maintaining low bit rates.
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining [16]. Please refer to quantitative results in  Sec. 4.1 for details.
Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
[…]
SwinIR [48](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD [92](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT [17])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT [18](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT [15](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L [15](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
ATD-F (Ours) 45.46 149.87 8,674 33.29 0.9082 29.48 0.8017 28.03 0.7539 28.53 0.8487 33.11 0.9335
DAT-F (Ours) 40.00 134.42 6,206 33.45 0.9094 29.60 0.8039 28.13 0.7560 28.75 0.8520 33.23 0.9339
HAT-F (Ours) 45.97 158.79 5,750 33.53 0.9100 29.65 0.8050 28.18 0.7569 28.79 0.8527 33.33 0.9342
HAT-F-ImageNet (Ours) 45.97 158.79 5,750 33.55 0.9102 29.63 0.8049 28.18 0.7569 28.80 0.8529 33.33 0.9342
HAT-L-F (Ours) 66.04 240.03 8,888 33.75 0.9116 29.87 0.8091 28.31 0.7597 29.51 0.8637 33.36 0.9343
HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
[…]
4 Experiments
4.1 Image Super-Resolution
Experimental Setup. We conduct extensive experiments to validate the effectiveness of our Factorized Features representation for Super-Resolution tasks. Following the strategy outlined in [15] and [20], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [37], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [22], which is used for [15] and [20] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [48], HAT [16], DAT [17], and ATD [92]) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K [1] + Flickr2K [52]) as the finetuning dataset following [21, 15] for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 2e-4, batch size 32, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Features, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
Quantitative Results. Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [6], Set14 [90], BSD100 [63], Urban100 [33], and Manga109 [64]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [91], DAT [17], and HAT [16]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts, ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [16] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [14, 16]. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods [82], the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in  Fig. 3. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison. We provide the visual comparison in Fig. 7. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Multi-frame SR. Due to space limitations, we provide additional comparisons on multi-frame SR in the supplementary materials.
(a) Kodak
(b) CLIC
(c) Tecnick
Figure 8: Performance (RD-Curve) evaluation on image compression using different datasets.
Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
Tot Enc $\downarrow$ Tot Dec $\downarrow$
VTM 0.00 129.21 0.14 -
Xie (MM’21) -0.78 2.93 6.00 50.0
Cheng (CVPR’22) 5.44 1.98 4.69 29.6
STF (CVPR’22) -4.31 0.14 0.13 99.9
ELIC (CVPR’22) -7.24 0.07 0.09 36.9
TCM (CVPR’23) -11.74 0.16 0.15 76.7
CCA (NeurIps’24) -20.97 0.11 0.09 64.9
TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
Method Kodak CLIC Tecnick
$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
[…]
TCM[55] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
4.2 Single- and Multi-Image Compressions
Experimental Setup. We evaluate our Factorized Features for image compression tasks against state-of-the-art methods. Following the Super-Resolution setup in Sec. 4.1, we train on the SA-1B dataset. To emphasize our representation, compression and decompression modules are initialized from pre-trained models [55], then fine-tuned end-to-end on 256×256 patches for 200k iterations using the AdamW optimizer [60] (learning rate=1e-5, batch size=16, betas=(0.9, 0.99)).
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 6.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
Rate-Distortion Performance Comparison. We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [56], [13], [98], [87], [19], [3], [44], [34], [66], [5], [71], and [32]. The classical image compression codec, VVC [80], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
In  Tab. 2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods. The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$. The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
Our experiments validate the effectiveness of FIPER in image compression, demonstrating substantial improvements across benchmarks (Kodak, CLIC, Tecnick). The multi-image compression strategy particularly excels with larger image sets, highlighting strong scalability. Additionally, our method achieves competitive latency, significantly improving compression efficiency while balancing quality and speed.
Table 4: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in Eq. 8.
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline [12] 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
4.3 Ablation Studies
Influence of SR Priors in Image Compression. We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness. TCM-HAT refers to using the original HAT [16] instead of our SR Module in Fig. 6b. TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly. Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Fields.
Effectiveness of Factorized Features Design. We conduct experiments to verify our multi-frequency modulation. The quantitative performance reported on single-image regression is shown in Tab. 4, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
5 Conclusion
We proposed Factorized Features, a representation that decomposes images into multi-frequency components to model implicit structures and patterns. Our approach addresses challenges in Super-Resolution and Image Compression by effectively restoring details and preserving visual fidelity. We integrate SR priors with Image Compression for better information recovery and introduce a basis-merging technique to enhance multi-image processing performance.
Limitations. Although effective, our method requires further optimization for computation-limited scenarios such as real-time decoding. Additionally, incorporating semantic information remains a promising direction for future research.
[…]
Supplementary Material
Appendix A Overview
This supplementary material provides additional implementation details and experimental results to complement our main manuscript. Specifically, we first elaborate on the background of factor fields and learned image compression in Appendix B, which were omitted in the main paper due to space limitations. Next, in Appendix D, we present comprehensive details about the benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper. Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in Appendix H.
Appendix B Preliminaries
In this section, we briefly introduce the background of the factor fields [12], and Learned Image Compression along with their properties.
Factor Fields.
The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently [11, 27, 68, 26, 7, 9], while [12] first proposed a unified representation called factor fields. For a 1D signal $s(x)$, the factor fields formulation is:
$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (11)
where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation. The coefficient and basis are queried by sampling with the coordinates $x$. This generalizes to Q-dimensional signals:
$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (12)
where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
Learned Image Compression.
Following [66, 3], a learned image compression model with a channel-wise entropy model can be formulated as:
$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (14)
$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},$
$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}),$
where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}})$. The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is then quantized to $\hat{z}$, where a factorized density model $\varphi$ is used to encode quantized $\hat{z}$ as $p_{\hat{z}\mid\varphi}(\hat{z}\mid\varphi)=\prod_{j}\left(p_{z_{j}\mid\varphi}(\varphi)*\mathcal{U}\left(-\frac{1}{2},\frac{1}{2}\right)\right)(\hat{z}_{j})$ where $j$ specifies the position or each element of each signal. The derivation of the equation can be found in [2]. Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
$L$ $=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x})$ (15)
$=\mathbb{E}\left[-\log_{2}\left(p_{\hat{{y}}|\hat{{z}}}\left(\hat{{y}}\mid\hat{{z}}\right)\right)\right]+\mathbb{E}\left[-\log_{2}\left(p_{\hat{{z}}|\varphi}\left(\hat{{z}}\mid\varphi\right)\right)\right]$
$+\lambda\cdot\mathcal{D}\left({x},\hat{{x}}\right)$
where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow [55], modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
Appendix C Differences from Factor Fields.
We explicitly highlight how sawtooth-based sampling and multi-frequency modulation extend the original factor fields. In our 1D and 2D toy examples as shown in Figs. 9 and 10, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$. Factor fields already improve accuracy over such baselines, but our introduction of $\psi$ and $\alpha$ further captures both high- and low-frequency details. Specifically, forcing $\alpha=4$ compels the network to learn broader frequency ranges and avoid noise from the fixed modulation. To confirm these advantages, we trained smaller SR models and compared them with factor fields in Tab. 5. The results illustrate that our multi-frequency constraint yields significant gains even in simpler configurations, underscoring its importance for high-frequency detail preservation and overall reconstruction quality.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Vanilla fields (b) Factor fields [10] (c) Ours with $\psi$ only (d) Ours with both $\psi$ and $\alpha$
Figure 9: 1D toy example.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Convergence curve (b) Target image (c) Vanilla fields (d) Factor fields (e) Ours with $\psi$ only (f) Ours with $\psi$ and $\alpha$
Figure 10: 2D toy example.
Table 5: Comparison with Factor Fields.
Method Sawtooth $\psi$ $\alpha$ Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
[…]
HAT-S 9.62 40.38 3,214 32.92 0.9047 29.15 0.7958 27.97 0.7505 27.87 0.8346 32.35 0.9283
HAT 20.77 86.02 3,692 33.04 0.9056 29.23 0.7973 28.00 0.7517 27.97 0.8368 32.48 0.9292
HAT-FA-Small 15.99 57.61 3,678 33.06 0.9065 29.26 0.7976 28.04 0.7524 28.12 0.8413 32.52 0.9301
HAT-F-Small-only-sin $\{\text{sin}\}$$\{1,4,16,64\}$ 16.13 66.34 4,176 33.16 0.9067 29.33 0.7989 28.05 0.7530 28.26 0.8429 32.64 0.9312
HAT-F-Small-half-freq $\{\text{sin},\text{cos}\}$$\{1,16\}$ 16.13 66.34 4,148 33.18 0.9073 29.35 0.8003 28.02 0.7521 28.31 0.8431 32.70 0.9316
HAT-F-Small $\{\text{sin},\text{cos}\}$$\{1,4,16,64\}$ 16.26 75.06 4,640 33.24 0.9079 29.39 0.8009 28.06 0.7543 28.41 0.8453 32.82 0.9324
HAT-F (Our FIPER) $\{\text{sin},\text{cos}\}$$\{1,4,16,64\}$ 45.97 158.79 5,750 33.31 0.9085 29.45 0.8011 28.09 0.7553 28.55 0.8463 32.99 0.9330
HAT-L (ImageNet pre-trained) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
Appendix D Implementation Details
Single Image Super-Resolution
Figure 11: Detailed architecture of Super-Resolution Modules
The architecture of our SR networks is shown in Fig. 11. Firstly, for Coefficient Backbones from various pre-trained models like [15, 91, 17], we replace the output upsampling layer (mostly pixel shuffle) with convolution of channel $D=256$, i.e., for input image with resolution $64\times 64$, the original pre-trained models process hidden states in $64\times 64$, and then upsample the output to $256\times 256$ for final result, where we replace the upsampling with convolution and output the final hidden state $X_{coeff}\in R^{256\times 64\times 64}$.
Next, the Basis Swin Transformer begins with feature embedding, extending the channels $D$ from $256$ to $384$, where the features are then passed through a series of Swinv2[59] and Sawtooth-Aware Downsampling Blocks. Specifically, each block contains two layers, one of which uses shifted window attention. After each block, the hidden state goes through upsampling by 4x and convolution to convert the output basis channel to $24$. Then, the Sawtooth-Aware Downsample block rearranges the hidden state $h$ as discussed in the main paper, which can be viewed as:
$h=\mathbf{rearrange}(h,\text{'b c (s h) (s w) $\rightarrow$ b (s s c) h w'}),\ s=2$ (16)
, and reduces the feature channel by 4. In all experiments we use $N=6$, i.e. there are 6 bases, where $Basis_{i}\in R^{24\times\frac{256}{2^{i}}\times\frac{256}{2^{i}}}\text{for i in }\{0\ ...\ 5\}$.
Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by Eq. 17:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)}.$ (17)
In practice, the coefficient is generated with channels equal to the total number of base channels, and it is split accordingly by channels.
Method Frames REDS4 Vimeo-90K-T Vid4
REDS/Vimeo PSNR SSIM PSNR SSIM PSNR SSIM
BasicVSR++[10] 30/14 32.39 0.9069 37.79 0.9500 27.79 0.8400
VRT[50] 16/7 32.19 0.9006 38.20 0.9530 27.93 0.8425
RVRT[49] 30/14 32.75 0.9113 38.12 0.9527 27.99 0.8462
PSRT[74] 16/14 32.72 0.9106 38.27 0.9536 28.07 0.8485
IA-RT[88] 16/7 32.90 0.9138 38.14 0.9528 28.26 0.8517
Ours-HAT-F 16/7 32.67 0.9096 38.09 0.9512 27.95 0.8441
Table 6: Performance comparison among different methods on REDS4, Vimeo-90K-T, and Vid4 datasets. Note that we don’t use any special design on the network, such as temporal correspondence or motion sampling like previous works, demonstrating our effectiveness.
Multi-Image (Video) Super-Resolution
Building upon our prior work in Single-Image Super-Resolution (SISR), we extend the applicability of Factorized Features to tackle the challenges in Multi-Image Super-Resolution (MISR). In the main paper, we introduced the Basis Merging Transformer for Multi-Image Compression. Here, we further expand its functionality to generate per-timestamp Coefficients and Bases.
To evaluate the effectiveness of our approach, we use Video Super-Resolution (VSR) as the benchmark task. Specifically, we simply leverage transformer blocks from [70] to model temporal correlations. For a given video, we first derive the Coefficient and Basis for each frame and rearrange each level $i$ of the Coefficients or Basis such that the spatial dimensions $H$ and $W$ are treated as the batch dimension, while the temporal dimension is processed as a sequence of tokens for the Transformer. Next, we apply positional embedding to the tokens, process the tokens with the transformer blocks, and finally decode the Coefficient and Basis token from the respective position with convolution layers for the output super-resolved frames.
Tab. 6 shows quantitative comparison with state-of-the-art methods. All VSR experiments were conducted using bicubic 4X downsampling. The training dataset includes the REDS[69] and Vimeo-90K[89] datasets, while the testing dataset comprices REDS4[69], Vid4[53], and Vimeo-90K-T[89].
For the REDS dataset, we train for 300k iterations using 16 input frames, with a learning rate of 1e-4 and a cosine learning rate that gradually decays to 1e-7. The batch size is set to 8, and the Coefficient Backbone and the Basis Swin Transformer are initialized from SISR. When training on the Vimeo-90K dataset, we first conduct 300k iterations with 14 input frames with flip sequence, we then train model on 7 input frames with flip sequence, using a learning rate of 1e-4 and a cosine learning rate decay to 1e-7. We initialize the weights using the model trained on the REDS dataset. The batch size remains 8. Note that the training is only conducted on the newly added transformer blocks, i.e., the Coefficient Backbone and Basis Swin Transformer are frozen. Test results for the REDS model are reported on the REDS4 dataset, while test results for the Vimeo-90K model are reported on Vimeo-90K-T and Vid4. We calculate PSNR and SSIM on the RGB channel for REDS4 and Y channel for Vimeo-90K-T and Vid4, following previous work[88, 74, 49].
Single Image Compression
Figure 12: Detailed architecture of Compression Modules
As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [55], as shown in Fig. 12(more details can be bound in [55]. What we do here is that we use the output from the last main layer and convert it to match the hidden channel of the SR Module.
Multi-Image Compression
Following VSR, we use transformer blocks to merge the bases. The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis. For the merged basis compression, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [55] with input and out channels change. After the Basis Merging Transformer is trained, we follow the same training setting for Basis compression while randomly sampling 1 to 24 images for a batch and setting only the Basis compression module trainable.
Appendix E Merged-Basis Strategy & SR Priors.
The SR prior accelerates detail recovery by leveraging high-resolution knowledge while basis merging exploits shared structures for efficiency. Since merging is introduced after, the compression model is pre-trained, and improvements stem not only from the SR prior but also from the well-initialized basis merging. When trained from scratch (see Tab. 7), the merging transformer tends to find a universal solution rather than image-specific structures. Additionally, merging inevitably causes information loss, prompting the coefficient branch to encode more image-specific details. This mutual offset between SR prior and basis merging ultimately strengthens our approach.
TCM-HAT-F-multi $M=24$ TCM-HAT-F-multi $M=24$ Coeff-Frozen TCM-HAT-F-multi $M=24$ Scratch
BD-rate % -20.97 -19.18 -18.34
Table 7: Comparison of Merged-Basis Strategies.
Appendix F Encoding/Decoding Times.
Fig. 5(c) processes each of the $M$ images through the Coefficient Backbone, Basis Swin Transformer, and Basis Merging Transformer, then compresses and transmits the merged basis. In Tab. 2, different $M$ values primarily affect the run time of the Basis Merging Transformer. Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in Tab. 8. Notably, Tab. 2 assumes each image’s basis is compressed independently, which can be confusing. In practice, a merged basis is processed only once, so averaged encoding/decoding time is more realistic.
$M=1$ $M=2$ $M=4$ $M=8$ $M=16$ $M=24$
Total Basis Merging Time (sec) 0.00328 0.00628 0.01140 0.02312 0.04784 0.08064
Averaged Basis Merging Time (sec) by $M$ 0.00328 0.00314 0.00285 0.00289 0.00299 0.00336
Averaged Merged Basis Compression Time (sec) by $M$ 0.0645 0.0323 0.0161 0.0081 0.0040 0.0027
Averaged Merged Basis Decompression Time (sec) by $M$ 0.0935 0.0467 0.0234 0.0117 0.0058 0.0039
Adjusted Enc Time (sec) by $M$ 0.2320 0.1998 0.1836 0.1756 0.1715 0.1702
Adjusted Dec Time (sec) by $M$ 0.1740 0.1272 0.1039 0.0922 0.0863 0.0844
[…]
Table 8: Comparison of Encoding/Decoding Times.
Appendix G Inference Time Comparison for SR.
We provide the inference time in Tab. 9, compared with extensive SR models. The inference time of different components of FIPER is also provided.
HAT-S HAT HAT-L ATD DAT RGT HAT-F-S (Ours) HAT-F(Ours) HAT-L-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
Inference Time (s) 0.0387 0.0398 0.0778 0.0592 0.0594 0.0612 0.0671 0.0716 0.1084 0.0400 0.0132 0.0184
Table 9: Inference time comparison for SR.
Appendix H More Visualization
H.1 Single Image Super-Resolution
Below, we provide more visual comparisons of single-image super-resolution, where FIPER denotes HAT-F.
[Uncaptioned image][Uncaptioned image]
H.2 Image Compression
Below, we provide more visual comparisons of image compression. The results are present with TCM-HAT-F and the metrics are PSNR/Bpp.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]